OpenStudy (anonymous):

Hello~Can someone help me with this math problem please and thanks in advance. find the indicated outputs for f(x)=2 to the 2nd power-5x f(0)= f(-1)= f(2)=

OpenStudy (mertsj):

\[f(x)=2^{2}-5x\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

almost 2x to the second power

OpenStudy (anonymous):

How did you do that

OpenStudy (mertsj):

I used the equation writer right below this box where it says equation.

OpenStudy (anonymous):

ohhhhhh thank you` I'm a newbie on here

OpenStudy (mertsj):

\[f(x)=2x ^{2}-5x\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So to find the f(0) replace x with 0. What do you get?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

oh wait for all of them? Let me try

OpenStudy (mertsj):

Yes. So f(0) = 0 and that is the ordered pair (0,0) Now find the f(-1) by replacing x with -1

OpenStudy (anonymous):

ohhhhhh- Thank you very much

OpenStudy (mertsj):

What did you get?

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

One moment-I'm still trying here :-)

OpenStudy (anonymous):

I got 20, but it is suppose to be an ordered pair right?

OpenStudy (anonymous):

(-4,5)?

OpenStudy (mertsj):

f(x)=2x^2-5x f(-1)=2(-1)^2-5(-1) f(-1)=2+5=7

OpenStudy (mertsj):

That would give us the ordered pair (-1,7)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

I'm going to look at your response to figure it out, I was way off

OpenStudy (mertsj):

Now find the f(2) by replacing x with 2: f(x)=2x^2-5x f(2)=2(2)^2-5(2) f(2)=8-10 f(2)=-2 So the ordered pair is (2,-2)

OpenStudy (anonymous):

for the first one how did you get the -7. I see how you got the -1. Oh wait you added the 5 and 2, correct?